OpenStudy (anonymous):

Find the rate of 5x+3y= -2

OpenStudy (anonymous):

If someone can explain this to me that would be great! I'll give a medal and become a fan who ever can help me (:

OpenStudy (anonymous):

5x+3y= -2 3y = -2 -5x y = -2/3 -5x/3

OpenStudy (anonymous):

thanks so much!!!!

OpenStudy (anonymous):

welcome @dusty :)